If there are enough rations for 15 troops for 8 days. Suddenly 5 more troops come into the group. How long would the rations now last. Thanks. It would be great to see how to set up the equation
Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52848)
Answer by Boreal(15235)

Number of troop-days is 15*8=120
With 20 then it is 20*x=120, x=6
another way is that days are directly proportional to number of rations and inversely proportional to number of troops
R(days)=k/T
8=k/15
k=120
now use k in the next part
R=120/20=6 days